OpenStudy (anonymous):

2(x + 1) = −8 Solve please! I am in such a anxious mood I can't even focus :C

OpenStudy (anonymous):

x= -5

OpenStudy (johnweldon1993):

You would distribute the 2 into the parenthesis 2x + 2 = -8 solve for x...

OpenStudy (anonymous):

Thanks. I am just ugh.

OpenStudy (johnweldon1993):

I hope everything's okay! just keep your head up! :)

OpenStudy (anonymous):

I got -3, I am not for sure if that is right, thanks I will try

OpenStudy (johnweldon1993):

Not quite 2x + 2 = -8 -2 -2 2x = -10 -------- 2 2 x = -5 :)

OpenStudy (anonymous):

Okie dokie!